Wilmington Company has two manufacturing departments—Assembly and Fabrication. It considers all of its manufacturing overhead costs to be fixed costs. The first set of data that is shown below is based on estimates from the beginning of the year. The second set of data relates to one particular job completed during the year—Job Bravo.

Estimated Data Assembly Fabrication Total
Manufacturing overhead costs $ 2,520,000 $ 2,880,000 $ 5,400,000
Direct labor-hours 90,000 54,000 144,000
Machine-hours 36,000 180,000 216,000
Job Bravo Assembly Fabrication Total
Direct labor-hours 19 11 30
Machine-hours 11 14 25

Required:

1. If Wilmington used a plantwide predetermined overhead rate based on direct labor-hours, how much manufacturing overhead would be applied to Job Bravo?

2. If Wilmington uses departmental predetermined overhead rates with direct labor-hours as the allocation base in Assembly and machine-hours as the allocation base in Fabrication, how much manufacturing overhead would be applied to Job Bravo?

(Round your intermediate calculation to 2 decimal places.)

1. Plantwide manufacturing overhead applied to Job Bravo

2. Manufacturing overhead applied from Assembly to Job Bravo

3. Manufacturing overhead applied from Fabrication to Job Bravo

4. Total departmental manufacturing overhead applied to Job Bravo
Answer #1

1) Plantwide manufacturing overhead rate = (5400000/144000) = 37.50

Plantwide manufacturing overhead applied to Job Bravo = 30*37.50 = 1125

2) Manufacturing overhead applied from Assembly to Job Bravo = 2520000/90000*19 = 532

3) Manufacturing overhead applied from Fabrication to Job Bravo = 2880000/180000*14 = 224

4) Total Departmental manufacturing overhead applied to Job Bravo = 532+224 = 756